# Tensor invariants of finite and classical groups

### v1

@@ 2.3. Finite generation
Let $S$ be a $\mathbf{GL}$-algebra and $R$ be a tca. The $\mathbf{GL}$-algebra $S$ is finitely generated in degrees $\leq s$ if it is generated (as a $k$-algebra) by the $\mathbf{GL}$-orbit of finitely many homogeneous elements all of degree $\leq s$. Similarly, the tca $R$ is finitely generated in degrees $\leq s$ if there exist finitely many elements $r_{i}$ in various $R_{n_{i}}$ with $n_{i}\leq s$, such that the smallest subspace of $R$ containing the $r_{i}$ and stable under multiplication and the action of the symmetric groups, is $R$ itself. The $\mathbf{GL}$-algebra $\operatorname{Sym}(U)$ (resp. tca $\operatorname{Sym}(\{V_{n}\})$) is finitely generated if and only if $U$ is a finite-length polynomial representation (resp. $\sum_{n}\dim(V_{n})$ is finite).
We also consider module objects for $S$ in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$; this forms an abelian category $\operatorname{Mod}_{S}$. The $\mathbf{GL}$-algebra $S$ is weakly noetherian if $S$-submodules of $S$ satisfy the ascending chain condition, and it is noetherian if $\operatorname{Mod}_{S}$ is locally noetherian. These definitions extend analogously to the tca $R$.
@@ 2.4. The Schur functor
$\mathcal{S}(V)\coloneqq\{\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},V)\}_{n\in\mathbf{N}}$
which can be identified with the $(1^{n})$ weight space of $V$ with $n$ varying (see [Rei16, § 2.4] and references therein). The Schur functor is an equivalence of symmetric monoidal categories if and only if $p=0$ (see [SS12, §5.4]).
@@
    The Schur functor $\mathcal{S}$ is exact;
@@
    For each partition $\lambda$, we have $\mathcal{S}(\mathbf{S}_{\lambda}(\mathbf{V}))=S^{\lambda}$, where $\mathbf{S}_{\lambda}(\mathbf{V})$ is the Schur module with highest weight $\lambda$ and $S^{\lambda}$ is the Specht module corresponding to $\lambda$.
@@
  1. (1)
    The exactness of $\mathcal{S}$ follows from, for example, the fact that $\mathbf{V}^{\otimes n}$ is a projective object in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$ (see [Rei16, Remark 2.10])
  2. (2)
    We have $\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},V)=(1^{n})$ weight space of $V$, and $(L_{\lambda}(\mathbf{V}))_{(1^{n})}\neq 0$ if and only if $\lambda$ is $p$-restricted (see [Rei16, § A.3]).
  3. (3)
    See [Rei16, § A.3]. ∎
@@
The Schur functor $\mathcal{S}$ is a symmetric monoidal functor.
@@
We first define a natural map $\mathcal{S}(U)\boxtimes\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$. Without loss of generality, we may assume $U$ and $W$ are finite length and homogeneous of degree $d$ and $e$ respectively. Recall that $\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},V)$ is naturally isomorphic to the $(1^{n})$ weight space of $V$ with the isomorphism given by $v\in V_{(1^{n})}\mapsto f_{v}$ where $f_{v}$ is the $\mathbf{GL}$-equivariant map which maps
$e_{1}\otimes e_{2}\otimes\ldots\otimes e_{n}\in\mathbf{V}^{\otimes n}\text{ to}v\in V.$
We define an $\mathfrak{S}_{d}\times\mathfrak{S}_{e}$-equivariant map from $\mathcal{S}(U)\times\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$ where $f_{u}\otimes f_{w}$ with $f_{u}\in\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes d},U)$ and $f_{w}\in\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes e},W)$ is mapped to $f_{u\otimes w}$. This extends to an $\mathfrak{S}_{d+e}$-equivariant map $\mathcal{S}(U)\boxtimes\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$ and is natural in both variables. Furthermore, the domain and codomain (for fixed $U$ and $W$) both have dimension $\binom{d+e}{d}\dim(U_{1^{d}})\dim(W_{1^{e}})$ and by comparing weight spaces, we see that the map is surjective, so it is an isomorphism. The natural isomorphism defined clearly commutes with the symmetric structures, so $\mathcal{S}$ is a symmetric monoidal functor. ∎
This proposition implies that the Schur functor takes a $\mathbf{GL}$-algebras and twisted commutative algebras. We will mainly use the following concrete application.
@@ 3. Finite group invariants
Throughout this section, we let $W$ be a representation of a group $G$, and $S$ and $R$ be the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})$ and tca $T(W)$ respectively.
@@ 3.1. The tca structure on $T(W)^{G}$
For finite groups $G$, the invariant subring $\operatorname{Sym}(W)^{G}$ is a polynomial ring only if the action of $G$ is generated by pseudo-reflections. Remarkably, Kharchenko [Kha78] proved that $T(W)^{G}$ is always free as an associative $k$-algebra. The tca $T(W)^{G}$ is almost never a free tca as can be seen using growth rates: $T(W)^{G}_{n}$ grows exponentially with $n$ because $\dim(T(W)^{G}_{n})\leq\dim(W)^{n}$ but a free tca with at least one generator of degree $\geq 2$ grows superexponentially. ∎
For the next result, we freely use the fact that the Schur functor $\mathcal{S}$ applied to an object $A$ can be identified with the subspace spanned by all vectors of weight $(1^{n})$ for all $n\geq 0$ (see Section 2.4). For us, a flat weight is a weight of the form $(1^{n})$ for some $n\geq 1$, and a flat weight vector is a weight vector of flat weight.
Lemma 3.2.
The Schur functor induces an identification
$\mathcal{S}(S^{G})\cong R^{G}.$
@@
We identify $R$ with the subspace of $S$ spanned by all flat weight vectors. Under this identification, the action of $G$ on $R$ is just the induced action on this subspace. Similarly, we identify $\mathcal{S}(S^{G})$ with the subspace of $S^{G}$ spanned by all flat weight vectors. An element is in $\mathcal{S}(S^{G})$ if and only if it (1) is a flat weight vector and (2) is also invariant under the action of $G$. By (1), the element is in $R$ and by (2), the element is in $R^{G}$. The reverse containment is similar. ∎
Corollary 3.3.
Assume $S^{G}$ is finitely generated in degrees $\leq r$. Then $R^{G}$ is also finitely generated in degrees $\leq r$.
Proof.
If $S^{G}$ is finitely generated in degree $\leq r$, there exists a finite length $\mathbf{GL}$-representation $U$ of degree $\leq r$ with a surjection
$\operatorname{Sym}(U)\to S^{G}.$
Applying the Schur functor, and using Lemma 2.3(1) and Lemma 3.2, we get a surjection
$\mathcal{S}(\operatorname{Sym}(U))\to R^{G}.$
By Proposition 2.4, we have $\mathcal{S}(\operatorname{Sym}(U))\cong\operatorname{Sym}(\mathcal{S}(U))$. Furthermore, the Schur functor $\mathcal{S}$ takes finite length objects of degree $\leq r$ to finite length objects of degree $\leq r$, we obtain the result. ∎
@@
Using Lemma 3.2 and the exactness of the Schur functor, we have that $R^{G}_{+}/(R^{G}_{+})^{2}\cong\mathcal{S}(S^{G}/(S^{G}_{+})^{2})$. By Lemma 2.3(ii), we get $\mathcal{S}(L_{\lambda}(\mathbf{V}))\neq 0$ for $p$-restricted $\lambda$. Therefore $R^{G}_{+}/(R^{G}_{+})^{2}$ is supported in infinitely many degrees, so $R^{G}$ is not finitely generated. ∎
When $p>0$, we do not know of an example where $R^{G}$ is finitely generated but $S^{G}$ is not. By the above result, all irreducible constituents of $S^{G}_{+}/(S^{G}_{+})^{2}$ in sufficiently large degree must not be $p$-restricted.
@@ 3.2. Finite groups
The key result that enables us to prove Theorem 1.1 is a mild refinement of Richman’s remarkable work [Ric96] about the failure of Weyl’s polarization theorem for finite groups in the modular setting. Richman shows that if $\operatorname{char}(k)\mid\#G$, then $S^{G}$ contains generators of sufficiently large degree – this result is insufficient for us as the Schur functor kills elements of large multidegree.
@@
Assume $G\subset GL(W)$ is a finite group with #$\#G$ divisible by $p$. For all $n\in\mathbf{N}$, there exists a $p$-restricted partition $\lambda$ with $|\lambda|>n$ such that $L_{\lambda}(\mathbf{V})$ is an irreducible constituent of $S^{G}_{+}/(S^{G}_{+})^{2}$.
@@
We explain how to obtain the result from [Ric96]. Set $m=n\frac{p^{\#G}-1}{p-1}$. In Proposition 9 of loc. cit., Richman shows that there exists a homogeneous element $f$ of degree $>\frac{m(p-1)}{p^{\#G}-1}=n$ which is nonzero in $S^{G}_{+}/(S^{G}_{+})^{2}$. This element $f$ contains a monomial $y$ that divides the monomial $y^{*}$ defined at the end of Page 36 of loc. cit.. The element $y^{*}$ is defined as the product $\prod(H_{i}(x_{i}))^{p-1}$ where $x_{i}$ is a linear form in $W$ and $H_{i}$ is the $i$-th inclusion $W\to W^{\oplus N}\subset W\otimes\mathbf{V}$. From this definition, we see that $y^{*}$ is a weight vector for $\mathbf{GL}$ of weight $\mu$ with $\mu_{i}=p-1$ for all $i$ with $\mu_{i}$ nonzero. So all nonzero components of the weight of $y$ is $<p$, and in turn, the same holds for $f$. Now, in a polynomial $\mathbf{GL}$-representation $U$ that has no $p$-restricted irreducible representation as a constituent, at least one component of each weight occurring in $U$ will be $\geq p$. Therefore, the $\mathbf{GL}$-representation generated by $f$ in $S^{G}/(S^{G}_{+})^{2}$ contains an irreducible $L_{\lambda}(\mathbf{V})$ with $\lambda$ $p$-restricted, as required. ∎
@@
The tca $T(W)$ is noetherian [SS17]. If #$\#G\neq 0$ in $k$, then $T(W)^{G}$ is a direct summand of $T(W)$ using the Reynolds operator. Therefore, $T(W)^{G}$ is also noetherian. To obtain the degree bound for $T(W)^{G}$, note that $\operatorname{Sym}(W\otimes\mathbf{V})^{G}$ is generated in degrees $\leq|G|$ by [Fle00], and therefore so is $T(W)^{G}$ by Corollary 3.3.
In the case when $p\mid\#G$, the invariant ring $T(W)^{G}$ is not finitely generated by Proposition 3.5 and Lemma 3.4. ∎
@@
Let $k$ be a field of characteristic $p>0$. Consider the $\mathfrak{S}_{p}$ action on $T(k^{p})=k\langle x_{1},x_{2},\ldots,x_{p}\rangle$. We identify $S=\operatorname{Sym}(k^{n}\otimes\mathbf{V})$ with $k[x_{i,j}|i\in[p],j\in\mathbf{N}]$. The invariant subring $S^{G}$ is not finitely generated [Ryd07]: the polynomials $f_{n}\coloneqq x_{1,1}x_{1,2}\ldots x_{1,n}+x_{2,1}x_{2,2}\ldots x_{2,n}+\ldots+x_{p,1}x_{p,2}\ldots x_{p,n}$ are nonzero in $S^{G}_{+}/(S^{G}_{+})^{2}$. Furthermore, up to scalars, the polynomials $f_{n}$ are the only nonzero elements of multidegree $(1^{n})$ in $S^{G}/(S^{G}_{+})^{2}$. The Schur functor maps $f_{n}$ to the element $p_{n}\coloneqq x_{1}^{n}+x_{2}^{n}+\ldots+x_{p}^{n}\in T(k^{p})$. So $\{p_{n}\}_{n}$ minimally generates $R^{G}$. ∎

### v2

@@ 2.4. Finite generation
Let $S$ be a $\mathbf{GL}$-algebra and $R$ be a tca. The $\mathbf{GL}$-algebra $S$ is finitely generated in degrees $\leq s$ if it is generated (as a $k$-algebra) by the $\mathbf{GL}$-orbit of finitely many homogeneous elements all of degree $\leq s$. Similarly, the tca $R$ is finitely generated in degrees $\leq s$ if there exist finitely many elements $r_{i}$ in various $R_{n_{i}}$ with $n_{i}\leq s$, such that the smallest subspace of $R$ containing the $r_{i}$ and stable under multiplication and the action of the symmetric groups is $R$ itself. The $\mathbf{GL}$-algebra $\operatorname{Sym}(U)$ (resp. tca $\operatorname{Sym}(\{V_{n}\})$) is finitely generated if and only if $U$ is a finite-length polynomial representation (resp. $\sum_{n}\dim(V_{n})$ is finite).
We also consider module objects for $S$ in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$; this forms an abelian category $\operatorname{Mod}_{S}$. The $\mathbf{GL}$-algebra $S$ is weakly noetherian if $S$-submodules of $S$ satisfy the ascending chain condition, and it is noetherian if $\operatorname{Mod}_{S}$ is locally noetherian. These definitions extend analogously to the tca $R$ (and in fact, for any commutative algebra object in a symmetric monoidal $k$-linear category).
@@ 2.5. The Schur functor
$\mathcal{S}(U)\coloneqq\{\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},U)\}_{n\in\mathbf{N}}$
which can be identified with the $(1^{n})$ weight space of $U$ with $n$ varying (see [Gre07, § 6.1]). The Schur functor is an equivalence of symmetric monoidal $k$-linear categories when $p=0$. When $p>0$, it is a $k$-linear functor that still satisfies many pleasant properties.
@@
    The Schur functor $\mathcal{S}$ is exact.
@@
    For all partitions $\lambda$, we have $\mathcal{S}(\mathbf{S}_{\lambda}(\mathbf{V}))=S^{\lambda}$.
@@
(1) is [Gre07, 6.2a]. (2) is obtained by combining [Gre07, 6.4a] and [Gre07, 6.4b]. (3) is [Gre07, 6.3(e)].
@@
The Schur functor $\mathcal{S}$ is symmetric monoidal.
@@
We first define a natural map $\varphi_{U,W}\colon\mathcal{S}(U)\boxtimes\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$. Without loss of generality, we may assume $U$ and $W$ are of finite length and homogeneous of degree $d$ and $e$, respectively. Recall that $\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},Z)$ is naturally isomorphic to the $(1^{n})$ weight space of $Z$ with the isomorphism given by $v\in Z_{(1^{n})}\mapsto f_{v}$, where $f_{v}$ is the $\mathbf{GL}$-equivariant map which maps
$e_{1}\otimes e_{2}\otimes\ldots\otimes e_{n}\in\mathbf{V}^{\otimes n}\text{ to }v\in Z.$
We define an $\mathfrak{S}_{d}\times\mathfrak{S}_{e}$-equivariant map from $\mathcal{S}(U)\times\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$ where $f_{u}\otimes f_{w}$ with $f_{u}\in\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes d},U)$ and $f_{w}\in\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes e},W)$ is mapped to $f_{u\otimes w}$. This extends to an $\mathfrak{S}_{d+e}$-equivariant map $\mathcal{S}(U)\boxtimes\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$ and is natural in both variables. Furthermore, the domain and codomain (for fixed $U$ and $W$) both have dimension $\binom{d+e}{d}\dim(U_{1^{d}})\dim(W_{1^{e}})$. By comparing weight spaces, we see that the map is surjective, so it is an isomorphism. This proves that the functor is monoidal, and $\varphi$ clearly commutes with the symmetric structures, so $\mathcal{S}$ is a symmetric monoidal functor as well. ∎
This proposition implies that the Schur functor takes $\mathbf{GL}$-algebras to twisted commutative algebras. We will mostly use the following concrete application.
@@ 3. Finite group invariants
Throughout this section, we let $W$ be a representation of a group $G$, and $S$ and $R$ be the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})$ and tca $T(W)$, respectively.
@@ 3.1. The tca structure on $T(W)^{G}$
For finite groups $G$, the invariant subring $\operatorname{Sym}(W)^{G}$ is a polynomial ring only if the action of $G$ is generated by pseudo-reflections. Remarkably, Kharčenko [Kha78] proved that $T(W)^{G}$ is always free as an associative $k$-algebra. However, the invariant ring $T(W)^{G}$ is never free as a tca unless $T(W)^{G}=T(W)$, as can be seen using growth rates: $T(W)^{G}_{n}$ grows exponentially with $n$ because $\dim(T(W)^{G}_{n})\leq\dim(W)^{n}$, but a free tca with at least one generator of degree $\geq 2$ grows superexponentially. ∎
For the next result, we freely use the fact that the Schur functor $\mathcal{S}$ applied to an object $A$ can be identified with the subspace of $A$ spanned by all vectors of weight $(1^{n})$ for all $n\geq 0$ (see Section 2.5). In what follows, a flat weight is a weight of the form $(1^{n})$ for some $n\geq 1$, and a flat weight vector is a weight vector of flat weight.
Proof of Theorem 1.2.
We identify $R$ with the subspace of $S$ spanned by all flat weight vectors. Under this identification, the action of $G$ on $R$ is the induced action on this subspace. Similarly, we identify $\mathcal{S}(S^{G})$ with the subspace of $S^{G}$ spanned by all flat weight vectors. An element is in $\mathcal{S}(S^{G})$ if and only if it (1) is a sum of flat weight vectors and (2) is also invariant under the action of $G$. By (1), the element is in $R$ and by (2), the element is in $R^{G}$. The reverse containment is similar.
For the second part, given $U\subset S^{G}$ that generates $S^{G}$, we obtain a surjection $\operatorname{Sym}(U)\to S^{G}$. Applying the Schur functor $\mathcal{S}$ and using Proposition 2.4 and Lemma 2.3(1), we get a surjection $\operatorname{Sym}(\mathcal{S}(U))\to\mathcal{S}(S^{G})\cong R^{G}$, as required. ∎
Corollary 3.2.
Assume $S^{G}$ is finitely generated as a $\mathbf{GL}$-algebra in degrees $\leq r$. Then $R^{G}$ is also finitely generated as a twisted commutative algebra in degrees $\leq r$.
@@
This easily follows from Theorem 1.2 and the fact that the Schur functor takes finite length objects to finite length objects. ∎
@@
By Theorem 1.2, we have an isomorphism $\mathcal{S}(S^{G}_{+})\cong R^{G}_{+}$. Furthermore, $(S^{G}_{+})^{2}$ is the image of the multiplication map $S^{G}_{+}\otimes S^{G}_{+}\to S^{G}_{+}$; the same holds for $(R^{G}_{+})^{2}$. Using the fact that $\mathcal{S}$ is a symmetric monoidal functor, we thus see that $\mathcal{S}((S^{G}_{+})^{2})\cong(R^{G}_{+})^{2}$, and by the exactness of $\mathcal{S}$, we also see that $R^{G}_{+}/(R^{G}_{+})^{2}\cong\mathcal{S}(S^{G}/(S^{G}_{+})^{2})$. By Lemma 2.3(ii), we have $\mathcal{S}(L_{\lambda}(\mathbf{V}))\neq 0$ for $p$-restricted $\lambda$. Therefore, if $S^{G}_{+}/(S^{G}_{+})^{2}$ has a non $p$-restricted component in arbitrarily large degrees, we have that $R^{G}_{+}/(R^{G}_{+})^{2}$ is also nonzero in infinitely many degrees or, $R^{G}$ is not finitely generated. ∎
When $p>0$, we do not know of an example where $R^{G}$ is finitely generated but $S^{G}$ is not. By the above result, all irreducible constituents of $S^{G}_{+}/(S^{G}_{+})^{2}$ of sufficiently large degree must not be $p$-restricted.
@@ 3.2. Finite groups
The key result that enables us to prove Theorem 1.1 is a mild refinement of Richman’s remarkable work [Ric96] regarding the failure of Weyl’s polarization theorem for finite groups in the modular setting. Richman shows that if $\operatorname{char}(k)\mid\#G$, then $S^{G}$ contains generators of sufficiently large degree – this result is insufficient for our purpose as the Schur functor kills elements of large multidegree. We begin with a preliminary result.
Lemma 3.4.
Assume $U$ is a polynomial representation of $\mathbf{GL}$ such that none of its irreducible constituents are $p$-restricted. If $w$ is a nonzero weight vector in $U$, the weight of $w$ has at least one component that is $\geq p$.
Proof.
It suffices to prove this result when $U$ itself is irreducible. Then, by the Steinberg tensor product theorem, we obtain an isomorphism $U\cong L_{\lambda}\otimes L_{\mu}^{(1)}$ where $\lambda$ is $p$-restricted and $\mu$ is nonempty, and $(-)^{(1)}$ is the functor that pulls back a representation along the Frobenius map $\mathbf{GL}\to\mathbf{GL}$. Every weight occurring in $U$ is therefore the sum of weights $\widetilde{\lambda}+p\widetilde{\mu}$ where $\widetilde{\lambda}$ is a weight occurring in $L_{\lambda}$, and similarly for $\widetilde{\mu}$. The result follows. ∎
The next result follows essentially from Richman’s work; we spell this out in the proof.
@@
Assume $G\subset GL(U)$ is a finite group with #G\#G divisible by $p$. For all $n\in\mathbf{N}$, there exists a $p$-restricted partition $\lambda$ with $|\lambda|>n$ such that $L_{\lambda}(\mathbf{V})$ is an irreducible constituent of $S^{G}_{+}/(S^{G}_{+})^{2}$.
@@
Choose $m\geq\frac{np^{\#G-1}-1}{p-1}$. By [Ric96, Proposition 9] there exists a nonzero homogeneous element in $S^{G}$, which we shall call $g$, of degree $\geq m\frac{p-1}{p^{\#G-1}-1}$ such that $g$ is nonzero in $S^{G}_{+}/(S^{G}_{+})^{2}$. We may assume that $g$ is a weight vector. In the proof of [Ric96, Proposition 7], Richman shows that the element $g$ can be chosen such that it contains a monomial that divides a nonzero monomial in the polynomial $f(W,m^{\prime})$ defined at the end of [Ric96, Page 30] for some $m^{\prime}\leq m$. The key point is that by definition, the polynomial $f(W,m^{\prime})$ is readily seen to be a weight vector for the $\mathbf{GL}$-action, with all nonzero components of its weight being p1p-1. Therefore, all components of the weight of $g$ are also <p<p. By the previous lemma, we get that the $\mathbf{GL}$-representation generated by $g$ in $S^{G}_{+}/(S^{G}_{+})^{2}$ must contain an irreducible representation that is $p$-restricted. ∎
@@
The tca $T(W)$ is noetherian [SS17]. If $G\neq 0$ in $k$, then $T(W)^{G}$ is a direct summand of $T(W)$ using the Reynolds operator. Therefore, $T(W)^{G}$ is also noetherian. To obtain the degree bound for $T(W)^{G}$, note that $\operatorname{Sym}(W\otimes\mathbf{V})^{G}$ is generated in degrees $\leq|G|$ by [Fle00], and therefore so is $T(W)^{G}$ by Corollary 3.2.
In the case when $p\mid\#G$, the invariant ring $T(W)^{G}$ is not finitely generated by combining Proposition 3.5 and Lemma 3.3. ∎
@@
Let $k$ be a field of characteristic $p>0$. Consider the $\mathfrak{S}_{p}$ action on $T(k^{p})=k\langle x_{1},x_{2},\ldots,x_{p}\rangle$. We identify $S=\operatorname{Sym}(k^{n}\otimes\mathbf{V})$ with $k[x_{i,j}|i\in[p],j\in\mathbf{N}]$. The invariant subring $S^{G}$ is not finitely generated [Ryd07]: the polynomials $f_{n}\coloneqq x_{1,1}x_{1,2}\ldots x_{1,n}+x_{2,1}x_{2,2}\ldots x_{2,n}+\ldots+x_{p,1}x_{p,2}\ldots x_{p,n}$ are nonzero in $S^{G}_{+}/(S^{G}_{+})^{2}$. Furthermore, up to scalars, the polynomials $f_{n}$ are the only nonzero elements of multidegree $(1^{n})$ in $S^{G}/(S^{G}_{+})^{2}$. The Schur functor maps the $\mathbf{GL}$-representation generated by the $f_{n}$ to the subspace generated by $p_{n}\coloneqq x_{1}^{n}+x_{2}^{n}+\ldots+x_{p}^{n}\in T(k^{p})$. So $\{p_{n}\}_{n}$ minimally generates $R^{G}$. ∎